# Decentralized COVID-19 Health System Leveraging Blockchain

### v1

@@ 3.2.3. Data User Module
The data user module is a functional module related to DU. Compared to DO, DU cannot upload data files. This module consists of four main parts: registration and login, key management, request sharing, and data query (Morris,, 2025). The function of the key management submodule is equivalent to the key management function introduced in DO. Due to DU’s inability to upload data, the data source for DU’s data queries comes entirely from request sharing, and the results that can be queried exist in the dataset that has received the sharing request.
@@ 4. System Design and Implementation Details
The system implementation part of this article is based on the Hyperledger Fabric architecture, as shown in Fig 3. The detailed system implementation can be roughly divided into three parts, namely the deployment of the Fabric network, SDK instantiation, and front-end design(Yu et al.,, 2024). The Fabric network mainly includes node definition and chain code installation. As the middle layer structure of the system, the SDK is responsible for deploying the system network downwards, conveying upper-layer operation instructions, and returning instruction execution results upwards. Front-end programs provide services through a web format.
@@ 4.1.1. Project Overview
Hyperledger (or Hyperledger Project) is an open-source project aimed at promoting the cross-industry application of blockchain. It was initiated by the Linux Foundation in December 2015, and its members include leaders in the finance, banking, IoT, supply chain, manufacturing, and technology industries. This project was created to promote further development of the blockchain industry and facilitate cooperation between various blockchain projects. This project will inherit independent open protocols and standards, using framework methods and dedicated modules, including consensus mechanisms and storage methods for each blockchain, as well as identity services, access control, and smart contracts (Mao et al.,, 2024).
Hyperledger Fabric is an excellent representative of consortium chains(Androulaki et al.,, 2018) and a platform that provides distributed ledger solutions. Each architecture in Fabric is modular and supports the combination of different components, allowing for the centralized processing of independent modules and the construction of an actual executable system.
According to the design concept of the system architecture, the Fabric architecture has proposed some new concepts, such as the concept of channel, which is an independent hyperledger fabric instance. This is the first concept related to blockchain systems proposed in Fabric. Channel is similar to the concept of a subnet in a network, where all channels are independent and do not depend on each other. Different channels have different rules and cannot share data with each other. The chain code mentioned earlier is closely connected to the channel during deployment, meaning that different channels need to deploy different chain codes (Zhang and Liew,, 2009).
@@ 4.1.2. Project Workflow
The Fabric blockchain defines three types of nodes, namely: Client (SDK), Peers, and Orderers. The client is similar to a client node and can make various requests to the system. The peers node records the ledger information in the blockchain system, while the orderers are similar to miners in the public chain, responsible for sorting and managing the update requests of the ledger and notifying the peers node to update the ledger. The three of them work together to form the entire blockchain network of Hyperledger Fabric(Foschini et al.,, 2020). The workflow of Fabric is shown in Fig. 4.
@@
In the update request operation, the SDK sends the read-write set results received from the Peer node to the Orderer node, requesting updates to the ledger. After receiving a certain amount of data, the Orderer node will process the data and first consider sorting the information. Its main purpose is to prevent the double-spending problem in the blockchain system. Then, it is necessary to check whether the digital signatures of the Peer nodes are correct and whether the simulated execution results of each Peer, that is, whether the read-write set is consistent, are checked. After all detections are passed, the Sorting node will send update instructions to all Peer nodes, and Peer nodes will update the ledger information based on the received content(Li et al., 2024b, ).
@@ 4.2.1. Channel Division
This is an advanced Fabric concept, as Fabric is a distributed system that typically involves multiple organizations, and different versions of Fabric code may (and typically) exist on different nodes within the network and channels within the network. Fabric allows this without requiring every peer node and sorting node to be at the same version level. In fact, supporting different version levels is the reason for supporting structural node rolling upgrades(Zhang et al.,, 2017).
@@ 14 Count: 1
An organization can contain multiple nodes and users, and a user can have multiple nodes. The above code defines an orderer node and an org organization containing two peers, with Domain identifying the respective domain names. Two Peer nodes belong to one user, and it is evident that one user does not possess decentralized characteristics (Li et al., 2024e, ). Therefore, this is only a rough definition for experimental testing.
@@
After the deployment of the relevant nodes is completed, a genesis block needs to be created, which is stored in the orderer node (Antal et al.,, 2021). Various information about the nodes needs to be written into the block to ensure the integrity of the data. All defined nodes are irreplaceable and tamper-proof. Once the data is uploaded, it is fixed, and others cannot impersonate a valid node or modify the content of the node, thus ensuring the security of the node data information and reducing the impact on the entire deployment. The creation of the genesis block is contained in the configtx.yaml folder. In addition to the information mentioned above being written into the block, it also defines some important parameter information about the blockchain system, including the sorting mode of sorting nodes, block generation time and capacity restrictions, specific code. The example is as follows:
@@ 4.2.3. Chain Code Design and Deployment
For the design process of a system based on the Hyperledger Fabric architecture, because the chain code is the only means of modifying the blockchain ledger, the design of the chain code is also the most important part (Li et al., 2024d, ). This is related to all the business functions that the designed system needs to complete, and it is also the key to distinguishing batches of different blockchain systems.
@@
Listing 4 shows the core code for updating data. Similarly, the number of input parameters for the function is checked first, and then the input data is deserialized (Liu et al., 2024a, ). Then check if the changed information is compliant, that is, the ID card information cannot be changed. After all checks are completed, assign all updated data values to complete the update operation.
@@
The general process of data query based on ID card number is also similar. The difference is that the function called by the query is more ID card-based and needs to query historical data. It calls the historical data query function stub. GetHistoryForKey under the development architecture (Li et al., 2024a, ).
@@ 4.3. System Network Service Startup
    Add system dependencies. The dependency information of the Go language is saved in the go.mod file. Enter the program folder and enter the go mod tidy command to add dependencies to the system (Wang et al.,, 2024).

### v2

@@ 3.2.3. Data User Module
The data user module is a functional module related to DU. Compared to DO, DU cannot upload data files. This module consists of four main parts: registration and login, key management, request sharing, and data query (Morris, 2025). The function of the key management submodule is equivalent to the key management function introduced in DO. Due to DU’s inability to upload data, the data source for DU’s data queries comes entirely from request sharing, and the results that can be queried exist in the dataset that has received the sharing request.
@@ 4. System Design and Implementation Details
The system implementation part of this article is based on the Hyperledger Fabric architecture, as shown in Fig 3. The detailed system implementation can be roughly divided into three parts, namely the deployment of the Fabric network, SDK instantiation, and front-end design(Yu et al., 2024). The Fabric network mainly includes node definition and chain code installation. As the middle layer structure of the system, the SDK is responsible for deploying the system network downwards, conveying upper-layer operation instructions, and returning instruction execution results upwards. Front-end programs provide services through a web format.
@@ 4.1.1. Project Overview
Hyperledger (or Hyperledger Project) is an open-source project aimed at promoting the cross-industry application of blockchain. It was initiated by the Linux Foundation in December 2015, and its members include leaders in the finance, banking, IoT, supply chain, manufacturing, and technology industries. This project was created to promote further development of the blockchain industry and facilitate cooperation between various blockchain projects. This project will inherit independent open protocols and standards, using framework methods and dedicated modules, including consensus mechanisms and storage methods for each blockchain, as well as identity services, access control, and smart contracts (Mao et al., 2024).
Hyperledger Fabric is an excellent representative of consortium chains(Androulaki et al., 2018) and a platform that provides distributed ledger solutions. Each architecture in Fabric is modular and supports the combination of different components, allowing for the centralized processing of independent modules and the construction of an actual executable system.
According to the design concept of the system architecture, the Fabric architecture has proposed some new concepts, such as the concept of channel, which is an independent hyperledger fabric instance. This is the first concept related to blockchain systems proposed in Fabric. Channel is similar to the concept of a subnet in a network, where all channels are independent and do not depend on each other. Different channels have different rules and cannot share data with each other. The chain code mentioned earlier is closely connected to the channel during deployment, meaning that different channels need to deploy different chain codes (Zhang and Liew, 2009).
@@ 4.1.2. Project Workflow
The Fabric blockchain defines three types of nodes, namely: Client (SDK), Peers, and Orderers. The client is similar to a client node and can make various requests to the system. The peers node records the ledger information in the blockchain system, while the orderers are similar to miners in the public chain, responsible for sorting and managing the update requests of the ledger and notifying the peers node to update the ledger. The three of them work together to form the entire blockchain network of Hyperledger Fabric(Foschini et al., 2020). The workflow of Fabric is shown in Fig. 4.
@@
In the update request operation, the SDK sends the read-write set results received from the Peer node to the Orderer node, requesting updates to the ledger. After receiving a certain amount of data, the Orderer node will process the data and first consider sorting the information. Its main purpose is to prevent the double-spending problem in the blockchain system. Then, it is necessary to check whether the digital signatures of the Peer nodes are correct and whether the simulated execution results of each Peer, that is, whether the read-write set is consistent, are checked. After all detections are passed, the Sorting node will send update instructions to all Peer nodes, and Peer nodes will update the ledger information based on the received content(Li et al., 2024b).
@@ 4.2.1. Channel Division
This is an advanced Fabric concept, as Fabric is a distributed system that typically involves multiple organizations, and different versions of Fabric code may (and typically) exist on different nodes within the network and channels within the network. Fabric allows this without requiring every peer node and sorting node to be at the same version level. In fact, supporting different version levels is the reason for supporting structural node rolling upgrades(Zhang et al., 2017).
@@ 14 Count: 1
An organization can contain multiple nodes and users, and a user can have multiple nodes. The above code defines an orderer node and an org organization containing two peers, with Domain identifying the respective domain names. Two Peer nodes belong to one user, and it is evident that one user does not possess decentralized characteristics. Therefore, this is only a rough definition for experimental testing.
@@
After the deployment of the relevant nodes is completed, a genesis block needs to be created, which is stored in the orderer node (Antal et al., 2021). Various information about the nodes needs to be written into the block to ensure the integrity of the data. All defined nodes are irreplaceable and tamper-proof. Once the data is uploaded, it is fixed, and others cannot impersonate a valid node or modify the content of the node, thus ensuring the security of the node data information and reducing the impact on the entire deployment. The creation of the genesis block is contained in the configtx.yaml folder. In addition to the information mentioned above being written into the block, it also defines some important parameter information about the blockchain system, including the sorting mode of sorting nodes, block generation time and capacity restrictions, specific code. The example is as follows:
@@ 4.2.3. Chain Code Design and Deployment
For the design process of a system based on the Hyperledger Fabric architecture, because the chain code is the only means of modifying the blockchain ledger, the design of the chain code is also the most important part. This is related to all the business functions that the designed system needs to complete, and it is also the key to distinguishing batches of different blockchain systems.
@@
Listing 4 shows the core code for updating data. Similarly, the number of input parameters for the function is checked first, and then the input data is deserialized (Liu et al., 2024). Then check if the changed information is compliant, that is, the ID card information cannot be changed. After all checks are completed, assign all updated data values to complete the update operation.
@@
The general process of data query based on ID card number is also similar. The difference is that the function called by the query is more ID card-based and needs to query historical data. It calls the historical data query function stub. GetHistoryForKey under the development architecture (Li et al., 2024a).
@@ 4.3. System Network Service Startup
    Add system dependencies. The dependency information of the Go language is saved in the go.mod file. Enter the program folder and enter the go mod tidy command to add dependencies to the system (Wang et al., 2024).